A single slit aperture has a width of 1.17 um.

A laser is directed at the slit, and the first observed minima (dark spot) occurs at an angle of 18.9 degrees.

What is the wavelength of the laser?
Answer #1

here,

width of slit , d = 1.17 um = 1.17 * 10^-6 m

theta = 18.9 degree

let the wavelength be lamda

we know

d * sin(theta) = n * lamda

for first order , n = 1

d * sin(theta) = lamda

1.17 * 10^-6 * sin(19.2) = lamda

lamda = 3.848 * 10^-7 m = 384.8 nm

the wavelength of laser light is 384.8 nm